​​​​​​MNM Corporation gives each of its employees an aptitude test. The scores on the test are normally distributed with a mean f 7
Answer #1

Solution :

Given that ,

mean = \mu = 75

standard deviation = \sigma = 15

n = 36

\mu\bar x = \mu = 75

\sigma\bar x = \sigma / \sqrt n = 15 / \sqrt 36 = 2.5

a) P(70.14 < \bar x < 82.14)  

= P[(70.14 - 75) / 2.5 < (\bar x - \mu \bar x) / \sigma \bar x < (82.14 - 75) / 2.5)]

= P(-1.944 < Z < 2.856)

= P(Z < 2.856) - P(Z < -1.944)

Using z table,  

= 0.9979 - 0.0259

= 0.9720

b) Using standard normal table,

P(Z \geq z) = 0.015

= 1 - P(Z \leq z) = 0.015  

= P(Z \leq z ) = 1 - 0.015

= P(Z \leq 2.17 ) = 0.985  

z = 2.17

Using z-score formula  

\bar x = z * \sigma \bar x+ \mu \bar x

\bar x = 2.17 * 2.5 + 75

\bar x = 80.425

C = 80.43

P(\bar x  \geq 80.43) = 0.015